OpenStudy (anonymous):

Solve 10^6x = 93. Round to the nearest ten-thousandth. Now the rounding part is not a problem. I tried solving this, but I never got one of the answer choices, which are A) 11.8109 B) 1.0986 C) 13.3801 D) 0.3281 Here's my logarithm: log10 of 93^6 = x

OpenStudy (mathmale):

\[10^{6x}=93\] Take either ln or log of both sides: \[\ln(10^{6x})=(\ln 93)\rightarrow 6x*\ln(10)=\ln 93.\]

OpenStudy (mathmale):

You'll need to evaluate those two logarithms. Now solve first for 6x, and then solve the resulting equation for x.

OpenStudy (mathmale):

this should lead you to the correct result.

OpenStudy (anonymous):

So the answer should be 0.3281?

OpenStudy (mathmale):

Perfect. Nice work!